# Toward Personalized LLM-Powered Agents: Foundations, Evaluation, and Future Directions

### v1

@@ 4. Memory
While profile modeling captures stable and high-level user traits, many fine-grained details essential for personalized responses cannot be fully encoded in a plain profile. To support such detailed personalization, agents rely on a memory module, which refers to the system’s ability to retain, recall, and leverage past interactions to improve future behavior (Weng, 2023). Memory is commonly divided into personal memory, which stores user inputs and interaction histories, and system memory, which records intermediate reasoning states or task-execution artifacts (Wu et al., 2025c). This survey focuses on personal memory, as it directly shapes how agents adapt to individual users. In the following, we first discuss how personal memory is structured and updated, and then how it is retrieved to support personalized behaviors.
@@
Personal memory can be implemented as internal memory, which embeds user-related information within the LLM’s internal components, or as external memory, which stores information in an auxiliary system. Internal memory approaches operate through modifications to model parameters (Ning et al., 2025a; Wang et al., 2025c; Jin et al., 2025a; Tan et al., 2024), key–value caches (Xi et al., 2024; Qian et al., 2025b), or hidden states (Zhang et al., 2025e, j; Wang et al., 2025e). Although parameter-efficient and compact (Wang et al., 2024e), internal memory is constrained by limited memory capacity, challenges in updating stored information, and reliance on model retraining, making it less suitable for dynamic, long-horizon personalization. By contrast, external memory stores information in an auxiliary system, usually combined with retrieval-augmented generation (RAG; Lewis et al., 2020), which enriches the input prompt with information retrieved from the external knowledge base, thereby improving user satisfaction and engagement. Compared to internal memory, external memory is more flexible and especially beneficial for users with limited personal data (Salemi and Zamani, 2025a). However, it requires careful design for storage, management, and personalized access. Given these considerations, we focus on external personal memory systems and structure this section around two key components: (1) personal memory schemes, which define how user-related information is stored and updated, and (2) personalized retrieval, which determines how stored information is accessed to guide personalized behaviors.
4.1. Personal Memory Schemes
Personal memory schemes define the foundation of the memory module, specifying how user-related information is organized, stored, and updated throughout interaction. Based on the timescale and influence of stored information, personal memory is often divided into short-term memory, which captures recent conversational context, and long-term memory, which accumulates enduring user information across sessions. Short-term memory provides reliable signals for immediate responses but is limited by context window constraints and cannot effectively utilize information beyond the current interaction session. On the other hand, long-term memory enables representation of long-tail individual data and supports self-evolution through accumulated experience (Jiang et al., 2024b), but suffers from information overload and difficulties in timely updating. As a result, recent personalized agents increasingly adopt hybrid memory designs that combine both short- and long-term memory to balance continuity and fine-grained adaptation (Li et al., 2024a), motivating careful design of the underlying memory structure and update mechanism.
@@ 4.1.1. Memory Structure
The memory structure determines what information is stored and how it is represented, forming the foundation for effective personalized retrieval. Early agent systems often appended full dialogue histories or execution trajectories to memory (Liu et al., 2024b), but such raw storage introduced redundancy, limiting the agent’s ability to surface user-relevant information. To achieve scalable personalized memory management, recent works convert interaction histories into compact and structured representations tailored to task demands (Liu et al., 2023; Zhong et al., 2024). Based on the format, memory structures can be broadly categorized as textual memory and structured memory.
@@
Textual memory stores user-related information in natural language form, typically by summarizing multi-turn interactions into concise and interpretable units. Because such representations preserve rich semantics and align naturally with the input–output format of LLMs, textual memory has become the dominant choice for personalized agents across diverse tasks (Zhang et al., 2025i). However, effective textual memory design faces two key challenges: defining appropriate segmentation boundaries and capturing comprehensive user-related information. Early approaches often relied on rigid turn-level or session-level segmentation, which frequently misaligned with the semantic units of dialogue and resulted in fragmented retrieval. To address this, recent work has explored more flexible, topic-coherent segmentation strategies. For instance, SeCom (Pan et al., 2025) and RMM (Tan et al., 2025c) partition conversations into semantically coherent segments, enabling memory units to reflect stable conversational topics rather than arbitrary boundaries. Nemori (Nan et al., 2025) further improves segmentation accuracy by using a boundary detector to group entire interactions into coherent episodes, preserving the integrity of user intent. Beyond segmentation, many systems enrich textual memory with auxiliary attributes such as timestamps, topic labels, and user personality (Liu et al., 2023; Xu et al., 2025d; memodb, 2025). For example, CAIM (Westhäußer et al., 2025a) introduces a tagging system that stores events with tuples of related tags, inductive thoughts, and timestamps. MemoryBank (Zhong et al., 2024) and PersonaAgent (Zhang et al., 2025c) adopt hierarchical summarization to model multi-level event structure and dynamic personality shifts, simulating the multi-component structure of human memory and enabling more coherent and human-like personalized interaction. MemInsight (Salama et al., 2025) is an autonomous memory augmentation approach that employs an attribute mining module to extract structured, semantically meaningful attributes from input dialogues. Several recent works draw inspiration from human memory systems (Shan et al., 2025; Zhang et al., 2025f). These systems typically distinguish between semantic memory and episodic memory (Liu et al., 2025c; Pink et al., 2025; Long et al., 2025). MIRIX (Wang and Chen, 2025) proposes an architecture with six specialized memory components (core, episodic, semantic, procedural, resource, and knowledge vault) organized hierarchically. Mem-$\alpha$ (Wang et al., 2025f) proposes an architecture with core, semantic, and episodic memory modules, combined with an RL framework that teaches agents to manage complex memory through interaction and feedback. MemWeaver (Yu et al., 2025) combines temporal and semantic dimensions across behavioral-level and cognitive-level memory to jointly represent user actions and long-term preferences. Collectively, these techniques advance textual memory toward richer representations, more coherent segmentation, and multi-level organization, allowing personalized agents to retain user-relevant details more faithfully and capture both explicit and implicit signals with higher accuracy.
@@
Structured memory provides a more explicit and computationally manipulable representation of interaction histories by storing information in predefined formats enriched with structural relationships. Compared with textual summaries, structured memory introduces stronger inductive biases that support scalable retrieval, multi-granularity reasoning, and more fine-grained personalization. Current approaches can be broadly grouped into vector-based memory, hierarchical tree structures, and graph-based memory architectures. Vector-based memory represents each memory unit using learned embeddings stored in a vector database, enabling scalable similarity search and fast retrieval of user-relevant information. This design has become a standard practice in personalized agents due to its efficiency and flexibility (Ocker et al., 2025; Chhikara et al., 2025; Wu et al., 2025d). However, vector memory alone discards structural relationships between memory items and offers limited interpretability, motivating more structured alternatives or hybrid systems that combine vectors with explicit organizational formats. Hierarchical tree structures organize memory at multiple abstraction levels, where each node corresponds to a memory unit and parent–child relationships encode increasingly coarse summaries. For instance, RAPTOR (Sarthi et al., 2024) recursively clusters and summarizes text into a tree, supporting retrieval at both fine- and coarse-grained levels. MemTree (Rezazadeh et al., 2024) similarly constructs a memory hierarchy in which each node stores aggregated textual content and semantic embeddings, allowing memory to be represented at varying abstraction depths. Graph-based architectures capture richer relationships by linking memory units through typed edges representing temporal, semantic, or relational dependencies (Jimenez Gutierrez et al., 2024; Li et al., 2025i; Chhikara et al., 2025). Some methods focus on behavioral coherence, such as MemWeaver (Yu et al., 2025), which builds an event-level behavior graph where interaction events are nodes connected by temporal and semantic edges, along with a context-aware random walk mechanism. Other designs integrate multi-dimensional subgraphs to capture more nuanced relationships between memory units. For example, AriGraph (Anokhin et al., 2024) combines semantic knowledge (entity–relation triplets) with episodic observation nodes, enabling both long-term knowledge accumulation and temporally grounded recall. Zep (Rasmussen et al., 2025) introduces a temporally-aware dynamic knowledge graph with three tiers, including episodic for non-lossy data storage, entity-level semantic for extracted entities and relationships, and community-level conceptual for high-level summarizations of strongly connected entity clusters. MemGAS (Xu et al., 2025a) further constructs cross-granularity associations across session-, turn-, summary-, and keyword-level memory to enrich personalized retrieval signals. While powerful, these graph-based structures often depend on computationally expensive LLM-based entity and relation extraction and may lose fine-grained contextual information during graph construction. To address this, SGMem (Wu et al., 2025d) proposes a lightweight sentence-graph memory, linking chunk nodes with their constituent sentences and connecting sentences via similarity edges to better preserve contextual semantics. Similarly, LiCoMemory (Huang et al., 2025g) introduces CogniGraph, a lightweight hierarchical indexing graph that organizes session summaries, entity–relation triples, and dialogue chunks through cross-layer hyperlinks, keeping the graph minimal while storing full content externally. Across these approaches, structured memory aims to provide richer organization by modeling how information relates, evolves, and aggregates across interactions.
@@ 4.1.2. Update mechanism
Memory update mechanism determines how an agent’s memory evolves as new interactions occur, ensuring that stored information remains accurate, consistent, and aligned with the user’s changing preferences. When a new memory unit arrives, the system must decide whether to incorporate it as a new entry, revise or merge existing ones, or prune outdated or contradictory content (Fang et al., 2025). Broadly, memory updating differs across two types of external memory: textual memory, where updates rewrite or reorganize summary-level information, and structured memory, where updates modify entity–relation structures through node/edge creation, conflict resolution, and temporal maintenance. We review the update mechanism for these two categories separately.
Textual memory stores interaction histories as natural-language segments or summaries. Updating therefore concerns how these segments are merged, rewritten, or reorganized when new dialogue arrives. Existing approaches can be grouped into similarity-driven updates and inference-guided updates.
@@
These methods update memory by comparing new segments with existing summaries and adjusting them based on semantic proximity and auxiliary metadata such as timestamps or topic labels. RMM (Tan et al., 2025c) summarizes each session by topic and updates the memory bank through semantic merging, replacing or consolidating summaries with high topical similarity. CAIM (Westhäußer et al., 2025a) extends this paradigm with a post-thinking step that generates structured, tag-annotated summaries from recent dialogue and integrates them into memory while refining related historical entries. Mem0 (Chhikara et al., 2025) introduces an LLM-driven update framework that chooses among ADD, UPDATE, DELETE, and NOOP operations for each new fact based on semantic similarity with existing memories. Memory-R1 (Yan et al., 2025a) builds on this with a reinforcement-learning memory manager that learns the optimal operation policy across multi-session interactions.
@@
Most personalized structured-memory systems update the memory by assessing semantic similarity between new information and existing units, then making local modifications to maintain coherence. In hierarchical tree structures, MemTree (Rezazadeh et al., 2024) updates memory by traversing the tree via semantic similarity, inserting new nodes through sibling attachment or leaf expansion, and incrementally updating ancestor nodes with LLM-based aggregation to preserve hierarchical coherence. Graph-based architectures demand even more complex updates, as every node may connect to multiple others (Huang et al., 2025g). $\text{Mem0}^{g}$ (Chhikara et al., 2025) embeds new triples, matches them to existing nodes, and establishes metadata-rich edges while using an LLM-based conflict resolver to deactivate outdated relations. Similarly, Zep (Rasmussen et al., 2025) integrates new episodic facts by timestamp alignment, updating or deactivating inconsistent edges and adjusting higher-level community clusters to maintain a temporally grounded semantic–episodic graph. MemGAS (Xu et al., 2025a) clusters new and historical multi-granularity chunks with GMM, selectively reinforcing only contextually relevant associations to ensure stable personalization over long horizons.
Reasoning- or agentic-driven updates
A smaller but emerging set of systems incorporate explicit reasoning, decision-making, or agentic inference to guide how memory should be updated, moving beyond similarity thresholds toward more cognitively aligned update dynamics. A-mem (Xu et al., 2025d) adopts an agentic update scheme in which each new memory becomes a structured note that triggers semantic linking and retroactive refinement of historical notes, enabling a continuously self-evolving memory graph. Although not designed for personalization, recent non-personalized graph-memory systems illustrate the potential of this direction. TRAIL (Zhao et al., 2025c) performs inference-driven updates by validating and refining candidate triples through confidence-guided decisions rather than pure similarity matching, while AriGraph (Anokhin et al., 2024) uses agentic reasoning to determine whether new observations should create links, update existing nodes, or refine episodic associations. These systems demonstrate that reasoning-based graph evolution can produce more coherent and cognitively grounded memory updates, suggesting a promising direction.
@@ 4.2. Personalized Retrieval
Memory retrieval determines which stored information becomes available to an agent at inference time and therefore directly influences the quality and consistency of personalized behaviors. Unlike conventional RAG systems, retrieval for personalized LLM-powered agents must simultaneously satisfy three requirements: (1) content relevance to the current query, (2) structural consistency with the underlying memory organization, and (3) personal alignment, ensuring that retrieved information reflects the user’s preferences. Existing methods address these challenges through three dominant mechanisms: content-based retrieval, structure-aware retrieval, and policy-guided retrieval.
@@
Content-based retrieval identifies relevant memory by measuring semantic or lexical similarity between the user’s query and stored information, typically implemented with dense retrievers (e.g., FAISS (Johnson et al., 2019), Contriever (Izacard et al., 2021)) or sparse lexical retrievers (e.g., BM25 (Robertson et al., 2009)). Retrieved segments are generally appended to the agent’s prompt to enrich contextual grounding (Wang and Chen, 2025; Kim et al., 2025c; Xu et al., 2025d). To capture comprehensive information about the user, content-based retrieval can be operated at multiple granularity levels. For example, RAPTOR (Sarthi et al., 2024) and MemTree (Rezazadeh et al., 2024) perform hierarchical content-based retrieval by selecting nodes whose embeddings are most similar to the query across different abstraction levels, even though the underlying tree structure is not explicitly traversed for reasoning. To further improve retrieval quality, some methods incorporate post-retrieval filtering or query refinement. Yuan et al. (2025) adopt a self-reflection mechanism that assesses whether the retrieved content is sufficient for response generation and refines the query when necessary. While effective at capturing semantic relevance, content-based approaches may overlook structural dependencies and long-range contextual relationships, motivating more structured retrieval designs.
@@
Structure-aware retrieval leverages the organization of memory, such as graph connectivity and episodic relationships, to retrieve information that is contextually linked to the query rather than merely similar in content. Such methods are often implemented in hierarchical or hybrid modes, where an initial retrieval of entities, topics, or relations guides subsequent retrieval of detailed memory units (Jimenez Gutierrez et al., 2024; Gutiérrez et al., 2025), potentially augmented with graph-based techniques such as Personalized PageRank (PPR; Bahmani et al., 2010). For instance, AriGraph (Anokhin et al., 2024) first retrieves semantically relevant relational triplets and then, conditioned on these triplets, retrieves the most relevant episodic memories. Zep (Rasmussen et al., 2025) combines semantic search with graph-structured signals, such as episode-mention frequency and node-distance scoring, to identify contextually relevant conversational episodes within its temporal memory graph. Other approaches enhance structure-aware retrieval with adaptive routing or self-reflection. MemGAS (Xu et al., 2025a) computes multi-granularity semantic similarities and integrates them using an entropy-based router, then applies PPR over a dynamically maintained association graph to retrieve structurally coherent memory units. AssoMem (Zhang et al., 2025b) performs associative retrieval by selecting clue nodes and linked utterances, followed by a multi-signal ranking procedure that fuses relevance, graph-based importance, and temporal alignment through a mutual-information–driven fusion strategy. By explicitly modeling the structural dependencies among memory units, structure-aware retrieval provides richer contextual grounding and improves the coverage of long-range or implicit user preferences.
@@
Policy-guided retrieval refers to retrieval strategies in which the selection of information is governed by user- or task-conditioned decision policies rather than static similarity matching (Zhong et al., 2024; Zhang et al., 2025c). Unlike content-based or structure-aware retrieval, which rely primarily on semantic proximity or memory topology, policy-guided retrieval actively determines how retrieval should proceed and can operate over arbitrary corpora without relying on a persistent user memory store. A first line of work focuses on pre-retrieval query transformation, where the retrieval intent is personalized before search. PBR (Zhang et al., 2025g) generates user-style pseudo feedback and corpus-anchored expansions to form a personalized query representation, while PersonaRAG (Zerhoudi and Granitzer, 2024) adapts document selection and ranking using real-time behavioral signals such as reading patterns and interaction context. A second line of work adopts policy-driven strategy selection, learning to invoke or integrate heterogeneous retrieval tools based on the task. UniMS-RAG (Wang et al., 2024b) formulates retrieval as a policy-learning problem, where supervised and reinforcement signals train the system to choose among keyword search, dense retrieval, and knowledge-graph lookup and to combine their results effectively. Persona-DB (Sun et al., 2025a) employs a two-stage persona-aware policy that first matches users via high-level persona keys and then combines retrievals from target users and collaborator users using a learned composition ratio to address sparsity and enhance personalization. A third direction incorporates schema-guided information completion, framing retrieval as a structured inference process. OCG-Agent (Shi et al., 2025b) decomposes narrative queries into an abstract data type schema and triggers a multi-route retrieval routine that fills schema fields in parallel; a reflect-and-complete loop then issues targeted supplementary retrievals until all required attributes are satisfied, followed by semantic normalization to unify heterogeneous results. Collectively, policy-guided retrieval shifts retrieval from passive document matching toward an active, decision-centric paradigm that adapts retrieval behavior to user preferences, task requirements, and interaction context, enabling more flexible and personalized information acquisition.

### v2

@@ 4. Memory
While profile modeling captures relatively stable and high-level user characteristics, personalized agents also require mechanisms for retaining user-related information across interactions. Memory provides this capability by enabling the system to store, recall, and leverage past experience to improve future behavior (Weng, 2023). Memory is commonly divided into personal memory, which stores user inputs and interaction histories, and system memory, which records intermediate reasoning states or task-execution artifacts (Wu et al., 2025d). This survey focuses on personal memory, as it directly shapes how agents adapt to individual users.
@@
Personal memory can be implemented either as internal memory, which embeds user-related information within the LLM, or as external memory, which stores such information in an auxiliary system. Internal memory may be realized through model parameters, key–value caches, or hidden states (Ning et al., 2025a; Wang et al., 2025c; Xi et al., 2024; Qian et al., 2025b; Zhang et al., 2025h; Wang et al., 2025e). Although compact and parameter-efficient (Wang et al., 2024d), internal memory is limited by constrained capacity, difficulties in updating stored information, and frequent reliance on retraining, making it less suitable for dynamic and long-horizon personalization. By contrast, external memory is more flexible and typically operates through retrieval-augmented generation (RAG; Lewis et al., 2020) to incorporate stored user information into the current context.
External personal memory can further be viewed at different temporal scales, including short-term memory, which retains recent conversational context, and long-term memory, which accumulates enduring user information across sessions. Short-term memory supports immediate responses but is constrained by context windows, whereas long-term memory enables the accumulation of long-tail personal information and continual adaptation (Jiang et al., 2024b), yet is more vulnerable to information overload and stale content. As a result, many personalized agents adopt hybrid designs that combine both. In the remainder of this section, however, we focus on long-term external personal memory, where the structured design of personalized memory has been most extensively studied. We discuss it from two perspectives: personal memory schemes, which determine how user-related information is stored and updated, and personalized retrieval, which determines how stored information is accessed to support personalized behavior.
4.1. Memory Schemes
Memory schemes determine how user-related information is organized, stored, and updated over time. In long-term external personal memory systems, this mainly involves the design of memory structure and update mechanism.
@@ 4.1.1. Memory Structure
Memory structure determines what information is stored and how it is represented, thereby shaping the scalability and usefulness of personal memory. Early agent systems often stored full dialogue histories or execution trajectories directly (Liu et al., 2024b), but such raw storage introduced redundancy and made it difficult to surface user-relevant information. Recent methods therefore transform interaction histories into more compact and task-adaptive representations (Liu et al., 2023; Zhong et al., 2024). Based on representation format, these structures can be broadly grouped into textual memory and structured memory.
@@
Textual memory stores user-related information in natural language form, typically by summarizing multi-turn interactions into concise and interpretable units. Because such representations preserve rich semantics and align naturally with the input–output format of LLMs, textual memory has become a common choice for personalized agents across diverse tasks (Zhang et al., 2025g). Its design mainly involves two key challenges: determining the granularity of memory units through appropriate segmentation, and preserving user-relevant information accurately within those units. Early approaches often relied on rigid turn-level or session-level segmentation, which could fragment semantically coherent interactions and hinder retrieval. Recent work therefore explores more flexible, topic-consistent segmentation strategies. For example, SeCom (Pan et al., 2025) and RMM (Tan et al., 2025c) partition conversations into semantically coherent segments, while Nemori (Nan et al., 2025) further improves segmentation by detecting episode boundaries and preserving the integrity of user intent. Once appropriate memory units are formed, a second challenge is to preserve user-relevant information faithfully and compactly. Many systems augment textual memory with auxiliary attributes such as timestamps, topic labels, and user personality (Liu et al., 2023; Xu et al., 2025d; memodb, 2025). For instance, CAIM (Westhäußer et al., 2025a) augments memory entries with tags, inductive thoughts, and timestamps, while MemInsight (Salama et al., 2025) extracts structured semantic attributes from dialogue for richer memory augmentation. More recently, some works draw inspiration from human memory systems by introducing multi-level organizations that distinguish semantic and episodic memory (Liu et al., 2025c; Pink et al., 2025; Long et al., 2025). For example, MIRIX (Wang and Chen, 2025) adopts a hierarchically organized multi-component memory architecture, and Mem-$\alpha$ (Wang et al., 2025f) combines core, semantic, and episodic memory with reinforcement learning for memory management. Overall, these efforts move textual memory toward more coherent segmentation, more faithful preservation of user-relevant content, and better support for long-term personalization.
@@
Structured memory represents interaction histories in predefined formats with explicit organizational relationships, making stored information easier to manipulate computationally. Compared with textual summaries, it introduces stronger inductive biases for scalable retrieval, multi-granularity reasoning, and fine-grained personalization. Current approaches can be broadly grouped into vector-based memory, hierarchical tree structures, and graph-based memory architectures.
@@
Vector-based memory encodes each memory unit as an embedding stored in a vector database, enabling efficient similarity search and fast retrieval of user-relevant information. Owing to its efficiency and flexibility, this design has become a common practice in memory modules (Ocker et al., 2025; Chhikara et al., 2025; Wu et al., 2025e). However, vector memory alone provides limited interpretability and does not explicitly capture relations among memory items, motivating more structured alternatives or hybrid systems that combine vectors with explicit organizational formats. Hierarchical tree structures organize memory at multiple abstraction levels, where parent–child relations encode increasingly coarse summaries. For example, RAPTOR (Sarthi et al., 2024) and MemTree (Rezazadeh et al., 2024) recursively cluster and summarize text into trees, supporting retrieval at both fine- and coarse-grained levels. Graph-based architectures capture richer relationships by linking memory units through typed edges that encode temporal, semantic, or relational dependencies (Jimenez Gutierrez et al., 2024; Li et al., 2025j; Chhikara et al., 2025). Some methods emphasize behavioral coherence. For example, MemWeaver (Yu et al., 2025) constructs an event-level behavior graph in which interaction events are connected by temporal and semantic edges, together with a context-aware random walk mechanism. Other designs instead integrate multi-dimensional subgraphs to model more nuanced relations among memory units. AriGraph (Anokhin et al., 2024) combines semantic knowledge with episodic observation nodes to support both long-term accumulation and temporally grounded recall, while Zep (Rasmussen et al., 2025) organizes memory into a temporally aware multi-tier knowledge graph spanning episodic, semantic, and conceptual levels. While expressive, graph-based memory often depends on costly LLM-based extraction and may lose fine-grained context during graph construction. Lightweight variants therefore simplify graph design while preserving contextual information. SGMem (Wu et al., 2025e) uses sentence-level graph organization, whereas LiCoMemory (Huang et al., 2025e) adopts a lightweight hierarchical indexing graph with full content stored externally. Overall, structured memory aims to provide richer organization by explicitly modeling how user-related information is organized, connected, and abstracted across interactions.
@@ 4.1.2. Update mechanism
The update mechanism determines how memory evolves as new interactions arrive, ensuring that stored information remains useful over time. When a new memory unit is observed, the system must decide whether to add it as a new entry, merge it with existing memory, revise outdated content, or discard irrelevant information (Fang et al., 2025). In external personal memory systems, updates differ between textual memory, where summary-level content is rewritten or reorganized, and structured memory, where nodes, edges, and their relations are modified. We discuss these two cases separately.
Textual memory stores interaction histories as natural-language segments or summaries. Updating therefore concerns how these units are merged, rewritten, or reorganized as new dialogue arrives. Existing approaches can be broadly grouped into similarity-driven updates and inference-guided updates.
@@
These methods update memory by comparing new content with existing entries and applying merge, revision, or replacement operations based on semantic similarity and auxiliary metadata. Earlier approaches, such as RMM (Tan et al., 2025c) and CAIM (Westhäußer et al., 2025a), perform summary-level merging and refinement. Mem0 (Chhikara et al., 2025) makes this process more explicit through ADD, UPDATE, DELETE, and NOOP operations over candidate memories, while Memory-R1 (Yan et al., 2025a) further learns the update policy with reinforcement learning across multi-session interactions.
@@
These methods update structured memory by matching new information to existing units through semantic similarity and then applying local structural modifications to preserve coherence. In hierarchical tree structures, MemTree (Rezazadeh et al., 2024) traverses the tree by semantic similarity, inserts new nodes through sibling attachment or leaf expansion, and incrementally updates ancestor summaries to maintain hierarchical consistency. In graph-based memory, updates are more complex because each node may connect to multiple others (Huang et al., 2025e). $\text{Mem0}^{g}$ (Chhikara et al., 2025) matches new triples to existing nodes, adds metadata-rich edges, and uses an LLM-based conflict resolver to deactivate outdated relations. Similarly, Zep (Rasmussen et al., 2025) updates episodic facts through timestamp-aware alignment while adjusting inconsistent edges and higher-level clusters, and MemGAS (Xu et al., 2025a) selectively reinforces contextually relevant cross-granularity associations to maintain long-horizon personalization.
Reasoning-guided updates
An emerging line of work uses explicit reasoning or decision-making to guide memory updates, moving beyond similarity matching toward more adaptive update dynamics. A-mem (Xu et al., 2025d) exemplifies this direction by treating each new memory as a structured note that triggers semantic linking and retroactive refinement of historical notes, enabling a self-evolving memory graph. Although not designed specifically for personalization, systems such as TRAIL (Zhao et al., 2025c) and AriGraph (Anokhin et al., 2024) further suggest that reasoning-guided graph evolution can support more coherent memory updates.
@@ 4.2. Personalized Retrieval
Memory retrieval determines which stored information becomes available to an agent at inference time and therefore directly influences the quality and consistency of personalized behaviors. Unlike conventional RAG systems, retrieval for personalized LLM-powered agents must simultaneously satisfy three requirements: (1) content relevance to the current query, (2) structural consistency with the underlying memory organization, and (3) personal alignment with the user’s preferences. Existing retrieval methods address these challenges through three dominant mechanisms: content-based retrieval, structure-aware retrieval, and policy-guided retrieval.
@@
Content-based retrieval selects relevant memory primarily through semantic or lexical similarity between the current query and stored information, typically using dense or sparse retrievers (Johnson et al., 2019; Izacard et al., 2021; Robertson et al., 2009). Retrieved content is then incorporated into the agent’s prompt to improve contextual grounding (Wang and Chen, 2025; Kim et al., 2025c; Xu et al., 2025d). This mechanism can also operate at multiple levels of granularity. For example, RAPTOR (Sarthi et al., 2024) and MemTree (Rezazadeh et al., 2024) retrieve memory across different abstraction levels based primarily on query similarity, even when the stored memory is hierarchically organized. Some methods further refine retrieval after the initial search. For instance, Yuan et al. (2025) uses self-reflection to assess whether the retrieved content is sufficient for response generation and revise the query when necessary. While effective for semantic matching, content-based retrieval does not explicitly model structural dependencies among memory items and may therefore overlook long-range contextual relations.
@@
Structure-aware retrieval accesses memory by explicitly exploiting graph, hierarchy, or episodic relations, rather than relying on content similarity alone. These methods often operate hierarchically, where the retrieval of entities, topics, or relations guides subsequent access to detailed memory units (Jimenez Gutierrez et al., 2024; Gutiérrez et al., 2025), sometimes with graph-based techniques such as Personalized PageRank (PPR; Bahmani et al., 2010). For example, AriGraph (Anokhin et al., 2024) retrieves relational triplets before retrieving relevant episodic memories, while Zep (Rasmussen et al., 2025) combines semantic search with graph-structured signals such as episode frequency and node distance. More adaptive variants include MemGAS (Xu et al., 2025a), which combines entropy-based routing with PPR over an association graph, and AssoMem (Zhang et al., 2025b), which performs multi-signal ranking over clue nodes and linked utterances. By modeling dependencies among memory units explicitly, structure-aware retrieval better captures long-range and implicit user preferences.
@@
Policy-guided retrieval refers to retrieval strategies in which memory access is controlled by user- or task-conditioned decision policies rather than by static similarity matching alone. Such a policy determines how retrieval should proceed, including how queries are formed, which retrieval tools are invoked, and whether additional information should be acquired. One line of work focuses on pre-retrieval query transformation, where personalized retrieval intent is constructed before search. For example, PBR (Zhang et al., 2025e) generates user-style pseudo feedback and corpus-anchored expansions to form a personalized query representation, while PersonaRAG (Zerhoudi and Granitzer, 2024) adapts document selection and ranking using real-time behavioral signals. A second line of work adopts policy-driven strategy selection, learning to invoke or combine heterogeneous retrieval tools according to task needs. UniMS-RAG (Wang et al., 2024b) formulates retrieval as a policy-learning problem over keyword search, dense retrieval, and knowledge-graph lookup, while Persona-DB (Sun et al., 2025a) employs a persona-aware policy to combine retrievals from target and collaborator users for improved personalization under sparse data. A third direction incorporates schema-guided information completion, framing retrieval as a structured inference process. OCG-Agent (Shi et al., 2025b) decomposes narrative queries into schema fields, retrieves relevant information through multiple routes, and iteratively supplements missing attributes until the schema is completed. Overall, policy-guided retrieval shifts retrieval from passive matching toward active decision making, enabling more flexible and personalized information acquisition.